# Dynamic Pricing with Adversarially-Censored Demands

### v1

@@ Appendix A Proof Details
When $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M-1,M$, we know that $r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})>0$ according to Lemma 5.4. Since $d_{t}(\frac{a+c-\gamma_{t}}{b})=a-b\cdot\frac{a+c-\gamma_{t}}{b}+N_{t}=\gamma_{t}-c+N_{t}\leq\gamma_{t}$ is not censored, we know that the optimal price $p_{t}^{*}$ satisfies $p_{t}^{*}>\frac{a+c-\gamma_{t}}{b}$ (since $r^{\prime}_{t}(p_{t}^{*})=0<r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})$) and therefore its demand is not censored. Therefore, the optimal price $p_{t}^{*}=\frac{a}{2b}$ and we have its regret bounded by Eq. 50 identically. Let $C_{non}:=\frac{2C_{s}(a_{\max}^{2}C_{b}^{2}+b_{\min}^{2}C_{a}^{2})}{4b_{\min}^{4}}$ and we have proven both cases. ∎

### v2

@@ 0.B.5 Proof of Lemma˜5
When $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M-1,M$, we know that $r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})>0$ according to Lemma˜3. Since $d_{t}(\frac{a+c-\gamma_{t}}{b})=a-b\cdot\frac{a+c-\gamma_{t}}{b}+N_{t}=\gamma_{t}-c+N_{t}\leq\gamma_{t}$ is not censored, we know that the optimal price $p_{t}^{*}$ satisfies $p_{t}^{*}>\frac{a+c-\gamma_{t}}{b}$ (since $r^{\prime}_{t}(p_{t}^{*})=0<r^{\prime}_{t}(\frac{a+c-\gamma_{t}}{b})$) and therefore its demand is not censored. Therefore, the optimal price $p_{t}^{*}=\frac{a}{2b}$ and we have its regret bounded by Eq.˜54 identically. Let $C_{non}:=\frac{2C_{s}(a_{\max}^{2}C_{b}^{2}+b_{\min}^{2}C_{a}^{2})}{4b_{\min}^{4}}$ and we have proven both cases.